OpenStudy (anonymous):

I need help solving the equation -y+6x=12

OpenStudy (dan815):

you cannot solve it completely

OpenStudy (dan815):

since you have a 2 variables in 1 equation

OpenStudy (dan815):

you can represent all its solutions in a graph though

OpenStudy (dan815):

re arrange in y=mx+b and graph this line

OpenStudy (dan815):

m is the slope, and b is the y intercept

OpenStudy (anonymous):

Yes that's correct I need to solve and graph

OpenStudy (dan815):

okay so lets rearrange in terms of y first so its easier to tink about

OpenStudy (anonymous):

ok

OpenStudy (dan815):

start by adding y to both sides what happens?

OpenStudy (anonymous):

I got rid of the y on left side

OpenStudy (dan815):

-y + y + 6x=12 + y these y cancel 6x=12+y can you see what you must do to isolate for y on the right side now

OpenStudy (anonymous):

do you write it like that

OpenStudy (dan815):

what do you mean?

OpenStudy (anonymous):

6x=12+y

OpenStudy (dan815):

oh yeah that is fine if it makes it easier to think u can flip them around since that is still the same equation like 12+y=6x

OpenStudy (anonymous):

ok

OpenStudy (dan815):

just think about it like someone is looking at the equation standing behind it, and u are looking at it standing infront of it, to you it might be on the right to the other guy its on his left, but they are still looking at the same equality right

OpenStudy (anonymous):

can you walk me through this problem

OpenStudy (dan815):

okay

OpenStudy (anonymous):

ty

OpenStudy (dan815):

so lets continue by isolating for y on 1 side all alone now

OpenStudy (anonymous):

ok

OpenStudy (dan815):

subtract 12 from both sides 12+y - 12=6x-12 y = 6x-12 ------------------ now we can graph this equation

OpenStudy (anonymous):

ok

OpenStudy (dan815):

to graph y=6x-12, there are many little tricks to do it fast, but for starters, you must realize this is the equation of a line since it is in the form y=mx+b where m is some random constant multiplying x

OpenStudy (anonymous):

ok

OpenStudy (dan815):

you can also infer it is a line by picking some x values and see what htey value must be there

OpenStudy (dan815):

we need atleast 2 points to graph, so lets try to find the simplest points by looking at when x=0, what is y and x=1, what is y

OpenStudy (dan815):

y=6x-12 when x=0 y=6*(0)-12 = -12, so we have our first point x=0 and y=-12 When x=1 y=6*(1)-12=-6, so we have our 2nd point x=1,y=-6

OpenStudy (anonymous):

ok im having a difficult time understanding this online thanks for your help but I need to talk to someone in person

OpenStudy (dan815):



OpenStudy (dan815):

okay wait I will give you online video to watch

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

thanks

OpenStudy (dan815):

on the left side there are other topics which you can browse by if you feel like you are missing something before you can understand graphing linear equations right away

OpenStudy (anonymous):

ty

OpenStudy (dan815):

okay, you are welcome! see you